Grade 12General Physics

where gravitational field is zero?

kanaram
10 Years agoGrade 12

1 Answer

Nikhil Upadhyay
10 Years ago
there are points between any two objects where gravititional field strength is zero
like in case of earth and moon there is a point which is 3.47 X 108 m from the centre of the earth where gravitational field intensity is zero.you can use this formula for finding the distance from a point where it is zero
E=F/M=GM/r2         (where F is the gravitational force)
if we put r=infinity  then E=0 therefore gravitational field decreases as distance r increases.